OpenStudy (peterg2001):

help will medal Subtract - 3x/11-5/11 Reduce, if possible.

OpenStudy (emmamink):

Ok, so it's \[\frac{ 3x }{ 11 }-\frac{ 5 }{ 11 }\]

OpenStudy (anonymous):

Not related to anything on here but @EmmaMink ur cat is adorable...is it yours?

OpenStudy (emmamink):

We can write this as \[\frac{ 3x-5 }{ 11 }\]

OpenStudy (emmamink):

No...sadly...but thanks Emma

OpenStudy (anonymous):

Your welcome..Its very cute :)

OpenStudy (emmamink):

I think that is as simplified as I can go @peterg2001

OpenStudy (emmamink):

Thanks for the medal!

OpenStudy (anonymous):

You did a great job explaining that @EmmaMink :) great job!